OpenStudy (anonymous):

Solve by factoring. 4x2 - 5x + 1 = 0

Parth (parthkohli):

\( \color{Black}{\Rightarrow 4x^2 - 4x - x + 1 = 0 }\) Does that help?

OpenStudy (anonymous):

maybe start by grouping: 4x(x-1)-(x-1)=0 then factor out (x-1): (x-1)(4x-1)=0...

OpenStudy (anonymous):

I'm working on Credit Recovery, and the work they're giving me; We've never worked on before.So, I don't really understand any of it.

OpenStudy (anonymous):